# To be or not to be, but where?

### v1

@@ 3.3 Emergence - Level 1: Classical Mechanics and Coarse Graining
Despite the change in the type of degrees of freedom, equations, and observables, the stage space remains physical space. To make this more explicitly, we can relax the condition of global thermal equilibrium and consider only local thermal equilibrium. In this case, the ideal gas law holds locally, and the temperature and pressure become scalar fields that vary from point to point in space. In this scenario, there are also dynamical equations, known as general transport equations, that explicitly incorporate time.
@@
Dynamical Equations $\vec{F}(\vec{x}_{\rm I})=m_{\rm I}\frac{d\vec{v}_{\rm I}}{dt}$ $PV=k_{\rm B}nT$
Table 2: Summary of the fundamental and emergent descriptions in the context of equilibrium thermodynamics. Once again, fundamentally there is a large number of degrees of freedom for individual particles (of order $n$ times Avogrado’s number in a room), while the emergent description simplifies the system by focusing on macroscopic thermodynamic variables like temperature, volume, and pressure. The emergent description retains the same stage space but significantly reduces the complexity by aggregating the behavior of numerous particles into a few macroscopic observables and a single equation.
@@ 3.4 Emergence - Level 2: Quantum Mechanics and Classical Mechanics
And about three decades later, quantum mechanics was developed, and the stage space itself would no longer be based on spacetime, but rather on quantum space, called Hilbert space.
Quantum mechanics also follows the structure of physical theories discussed above. Quantum mechanical degrees of freedom are typically represented as vectors in Hilbert space [^13] In this section, this simplification suffices to the points we want to make., evolving according to the Schrödinger equation. Physical observables are represented by self-adjoint operators instead of smooth measurable functions. However, at face value, the theory introduces other structures, such as the Born rule and the (real or apparent) collapse of the quantum state: the Schrödinger equation is linear, it allows for superpositions of different states, but it seems that measurements unavoidably reduce the superimposed quantum state to a single one. In other words, measurements seemingly give out single outcomes. This is also known as the measurement problem [^14] Given the extensive literature, among personal favorites are [16, 17, 18, 19] and will be further discussed in Sec. 8.
Now, it is generally believed that the world is fundamentally quantum mechanical. After all, it rules the small, microscopic world, and the macroscopic world is composed of these small entitites, so it seems that classical mechanics has lost its status as a base theory, and should be seen as the emergent description of the macroscopic world instead. As this time around the state space also changes, we call this Emergence - Level 2. We summarize the example of the classical emergent description from quantum mechanics in Table 3.
@@
Dynamical Equations $i\hbar\frac{d}{dt}|\psi\rangle=\hat{H}|\psi\rangle$ $\vec{F}(\vec{x}_{\rm I})=m_{\rm I}\frac{d\vec{v}_{\rm I}}{dt}$
@@ 4 To be or not to be, but where?
This is evident from our language. In fact, ontology, understood as the study of what there is and their relations [20], relies on the verb to be for its definition. This verb is inherently intransitive and can stand alone to define existence. However, it reflects the physical intuition that entities exist in space and time, as no other type of stage space was considered until the advent of quantum mechanics, at least in physics [^15] In the philosophy of mind, non-physicalist philosophers have introduced mental states that do not have the material world as their stage space. The ontological challenges of reconciling the classical and quantum stage spaces can draw parallels from that literature, as explored in [21, 22], where the hard problem of spacetime is discussed.. Thus, we are now faced with two possibilities: i) to move from a classical ontology (where the verb to be is defined in spacetime) to a quantum ontology (where the verb to be is defined in Hilbert space), or ii) to allow for a multi-level ontology [11], reassured by the notion of classical real patterns emerging from the underlying quantum world [13].
@@ 4.2 What are the possibilities?
Another important aspect of this discussion is the expected plurality of entities. If we argue for the universe as a whole, either classically or quantum-mechanically, we talk about things in the singular and focus on the holistic aspects of either spacetime or Hilbert space, their global state, and Hamiltonian as what ultimately exists. The partition of these spaces into what we typically see as classical and quantum subsystems is viewed as emergent, despite the persistence of real patterns associated with emergent entities [14, 13]. From this perspective, the second part of the ontological project becomes superfluous, as there is no meaningful notion of relations for a set composed of a single element, except the trivial comparison with the empty set [^17] An instance of this comparison in physics can lead to questions such as Why is there something rather than nothing? [25].. A recent defense of this position can be found in [26]. This holistic approach will be contrasted with the atomic one where there exists more than one single irreducible entity and relationships among them.
@@ 4.4 Quantum Ontology
The context of quantum mechanics makes everything much more interesting. In quantum mechanics, the primary elements of reality are not particles localized in spacetime but quantum states described by vectors in Hilbert space. These states can exist in superpositions and exhibit spatiotemporal non-local correlations, challenging the classical notion of independence. Adopting a quantum ontology thus requires rethinking the independence and interaction of parts. This perspective forces a reconsideration of how we define and measure the components of a system, as well as how we understand their interactions—topics we will explore in Sec. 7 and Sec. 8. Quantum mechanics seems to suggest a more holistic view, where the properties of the whole system cannot always be decomposed into the properties of its parts. This reasoning also justifies approaches to quantum cosmology and views where reality is seen as an evolving vector in Hilbert space [26].
@@
From my perspective, the main reason to consider this ontology is to attribute a sense of reality to entities that satisfy the Schrödinger equation, in contrast to quantum subsystems that seemingly demand some form of non-unitary evolution, apparent or not. Thus, the measurement problem will play a crucial role in how we approach this topic in Sec. 8.
@@
explores the possibility of maintaining a form of atomism within the framework of quantum mechanics, where the universe is composed of distinct quantum subsystems. Unlike classical atomism, which relies on spatial localization and independence, each quantum subsystem is associated either with factors in Hilbert space for finite-dimensional systems (Sec. 2) or via local operator algebras in algebraic quantum field theory (Sec. 5). However, the existence of entanglement undermines the statistical independence of quantum systems, at least as long as they are understood as invariant under the unitary evolution of the global Hilbert space to which they belong. We will revisit this in Sec. 8, where we will offer an alternative way to bridge holistic and atomistic quantum ontology, providing a new perspective on the measurement problem.
@@ 4.5 How do we decide?
As mentioned briefly above, we seem to be trapped in spacetime with classical measurements. Thus, given our invocation of scientific realism, which employs rationales such as the no-miracles argument [27] to attribute a sense of reality to theoretical terms in our theories, we soon are led into problems, such as our mysterious box in Sec. 1. There, we saw that under our limited measurements—only the energy displayed by the box—we could not infer the nature of the quantum system inside the box.
As discussed in Sec. 2, atomistic classical and quantum ontologies are intrinsically interconnected by the requirement that local measurements can be performed independently, establishing a direct connection between macro- and microcausality. Now, if there is ambiguity at the quantum level about how the Hilbert space is partioned and how local the quantum subsystems are interacting, we could imagine that further criteria should be imposed to select a preferred factorization of the Hilbert space, or assignment of local algebras. This project is pursued to some extent [^19] Other ideas for selecting a preferred factorization of the Hilbert space have been recently pursued in [6, 7, e.g.], with general discussions about decompositions of Hilbert space found in [32]. in [5] where they argue that imposing a quantum mereology resulting in the most possible quantum-local Hamiltonian is (almost always) sufficient to select a particular TPS, which then leads to a specific classical limit. From the perspective of classical ontology, we can invert this picture and view their program (which could be thought of as establishing the grounds for discussing quantum mereology from a Hilbert space fundamentalist perspective) as recognizing that we indeed live in a specific classical limit and then finding the necessary criteria at the quantum level aligned with this classical limit.
@@ 5 Status-of-art, or crash course on algebraic quantum field theory
At the end of the day, we are seemingly trapped in spacetime. By this, I mean that the real patterns [14] surrounding us are best described by spacetime physics, which includes all the possible measurements we make [33]. In fact, despite their complexity, measurements still amount to simply counting—such as the number of clicks in a detector or spots on a screen—from which we can infer the state of our physical systems.
@@ 5.1 What could we have hoped for?
We operate under the belief that the world is fundamentally quantum-mechanical and that the entire spacetime is associated with an underlying global Hilbert space. Here we see the relevance, and direct connection between the holistic ontologies discussed in Sec. 4.2. A crucial aspect of our activities in spacetime is the idea that it can be divided into independent parts. This concept underlies the separation between a system to be measured, our devices, and even ourselves as observers. The question then becomes what is needed to ensure that the whole can be broken into independent pieces, and to what extent this can be realized in Hilbert space.
Unexpectedly, this is a deep philosophical and physical problem that also appears in many other disciplines. We do not intend to provide an exhaustive discussion of the matter, but we will highlight at least one criterion that we already met that is needed classically to define any reasonable notion of independence: causal independence. If we know that two systems cannot exchange any information in a way that all possible observers would agree on, this is a strong criterion for considering them as independent entities. Special and general relativity provide the framework where this can be formulated. We will refer to this causal independence as spatiotemporal locality, as it establishes the condition under which we can securely say that something is independently located relative to something else (if they were not causally disconnected, there would be observers for whom the systems could be arbitrarily close to each other).
Let’s be more precise. What we have in mind are two small regions of spacetime, $U$ and $V$, which are spacelike separated and define local Cauchy surfaces. This means that there is a spacelike curve inside each region that contains all the necessary information to describe anything inside those regions.
@@
If we recall our initial mysterious box, perhaps this should not come as a surprise. Although quite often locality in Hilbert space, in the sense of local subsystems with commuting algebras, is a direct consequence of spatiotemporal locality (as discussed in Sec. 2), this is mostly our spacetime parochialism. As emphasized in recent works [34, 5, 6, 7, e.g.], the notion of spatiotemporal locality should be seen as emergent from the quantum formalism. So what can be done instead?
@@ 5.2 The Algebraic Approach
By analogy with coarse-grained emergent descriptions discussed in Sec. 3.2 and 3.3, in the space of quantum models aimed at reproducing classical spacetime measurements, is there a notion of an ‘ensemble’ of models that all converge to the same set of numbers one writes down in the lab? Is there a notion of model coarse-graining that needs to be considered because, similar to our accessible macrostates in thermodynamics, we only have access to classical measurements in the end? We will return to these questions in Sec. 8, but we have already seen that in the case of dualities, this is exactly the situation: different sets of commuting subalgebras can be used to arrive at the same classical measurements, even though they give rise to different spatiotemporal local structures, as observed in Sec. 2.
This highlights the subtleties involved in defining quantum subsystems and the emergent nature of spatiotemporal locality from a more fundamental quantum framework. And now, at least we finally understand why the maps between the different sets of spin sites were non-locally related since the choice of commuting subalgebras is one of the criteria that leads to the notion of independent subsystems.
@@ 5.4 More ambiguities!
If we could turn off the gauge interaction, or effectively deactivate the scalar field charge, we could consider the scalar field itself as a local observable (assuming some smearing) associated with different spacetime regions, such that for any two spacelike-separated regions $x\in U$ and $y\in V$, microcausality holds:
@@
where $\Gamma$ is a curve connecting $x$ and $x^{\prime}$ and it represents a Wilson line connecting two filed excitations of opposite charges. Then, as long as $D(x,x^{\prime})$ can be localized in $U$, it will commute with analogous operators defined in other spacelike-separated regions, and once again we find a gauge-invariant algebra of observables that is microcausal.
@@ 6 What about gravity?
Any good story starts with a mystery and at some point introduces a villain. Our understanding of gravity has dramatically changed since Newton’s theory. Initially, space and time were deemed absolute, and gravity was yet another force acting on this absolute stage. However, we have since learned with Einstein that the stage itself is dynamical, changing depending on the matter-energy act at play, and this very dynamics is responsible for gravitational phenomena.
@@
Thus, it is quite clear that Einstein’s perspective does not take space, and perhaps time, as independent substances. However, Einstein’s general relativity did not quite perfectly achieve this strict version of relationalism as it does allow vacuum solutions, but it was a significant step in that direction. Regardless, general relativity is extremely successful and underlies both astrophysics and modern cosmology, each presenting a large body of evidence towards its validity across different physical regimes.
@@
Nonetheless, many lessons have been learned so far. But we will focus on one lesson that I take to be underappreciated, and will lead us to reconsider several approaches to important problems in different areas of physics in Sec. 7. This is the problem of defining local gauge-invariant observables in linearized quantum gravity.
@@ 6.1 Operational Measurement Limitation
Let’s start intuitively [46, see interview with Nima Arkani-Hamed]. Imagine you are back in your lab, and you want to make a perfectly sharp measurement, or at least as sharp as possible. The problem is how to store that information in your apparatus. Quantum mechanically, you could imagine having an infinite chain of qubits to record any measurement using binary digits. This would take a lot of space, so perhaps you have access to a quantum field that has an infinite-dimensional Hilbert space at each point in space. Either way, you need an infinite-dimensional Hilbert space for a perfectly sharp measurement. While this may not be feasible in practice, there is nothing in quantum mechanics that prevents you from imagining it in principle.
@@
Now, let’s explore more concretely what happens when we try to define local gauge-invariants observables in gravity.

### v2

@@ 3.3 Emergence - Level 1: Classical Mechanics and Coarse Graining
Despite the change in the type of degrees of freedom, equations, and observables, the stage space remains physical space. To make this more explicit, we can relax the condition of global thermal equilibrium and consider only local thermal equilibrium. In this case, the ideal gas law holds locally, and the temperature and pressure become scalar fields that vary from point to point in space. In this scenario, there are also dynamical equations, known as general transport equations, that explicitly incorporate time.
@@
Dynamical Equations $\vec{F}_{I}(\vec{x}_{\rm I})=m_{\rm I}\frac{d\vec{v}_{\rm I}}{dt}$ $PV=k_{\rm B}nT$
Table 2: Summary of the fundamental and emergent descriptions in the context of equilibrium thermodynamics. Once again, fundamentally there is a large number of degrees of freedom for individual particles (of order $n$ times Avogadro’s number in a room), while the emergent description simplifies the system by focusing on macroscopic thermodynamic variables like temperature, volume, and pressure. The emergent description retains the same stage space but significantly reduces the complexity by aggregating the behavior of numerous particles into a few macroscopic observables and a single equation.
@@ 3.4 Emergence - Level 2: Quantum Mechanics and Classical Mechanics
And about three decades later, quantum mechanics was developed, and the stage space itself would no longer be based on spacetime, but rather on quantum space: the Hilbert space.
Quantum mechanics also follows the structure of physical theories discussed above. Quantum mechanical degrees of freedom are typically represented as vectors in Hilbert space [^16] In this section, this simplification suffices for the points we want to make., evolving according to the Schrödinger equation. Physical observables are represented by self-adjoint operators instead of smooth measurable functions. However, at face value, the theory introduces other structures, such as the Born rule and the (real or apparent) collapse of the quantum state: the Schrödinger equation is linear, it allows for superpositions of different states, but it seems that measurements unavoidably reduce the superimposed quantum state to a single one. In other words, measurements seemingly give out single outcomes. This is also known as the measurement problem [^17] Given the extensive literature, among personal favorites are [21, 22, 23, 24] and will be further discussed in Sec. 8.
Now, it is generally believed that the world is fundamentally quantum mechanical. After all, it rules the small, microscopic world, and the macroscopic world is composed of these small entities, so it seems that classical mechanics has lost its status as a base theory, and should be seen as the emergent description of the macroscopic world instead. As this time around the state space also changes, we call this Emergence - Level 2. We summarize the example of the classical emergent description from quantum mechanics in Table 3.
@@
Dynamical Equations $i\hbar\frac{d}{dt}|\psi\rangle=\hat{H}|\psi\rangle$ $\vec{F}_{I}(\vec{x}_{\rm I})=m_{\rm I}\frac{d\vec{v}_{\rm I}}{dt}$
@@ 4 To be or not to be, but where?
This is evident from our language. In fact, ontology, understood as the study of what there are and their relations [25], relies on the verb to be for its definition. This verb is inherently intransitive and can stand alone to define existence. However, it reflects the physical intuition that entities exist in space and time, as no other type of stage space was considered until the advent of quantum mechanics, at least in physics [^18] In the philosophy of mind, non-physicalist philosophers have introduced mental states that do not have the material world as their stage space. The ontological challenges of reconciling the classical and quantum stage spaces can draw parallels from that literature, as explored in [26, 27], where the hard problem of spacetime is discussed.. Thus, we are now faced with two possibilities: i) to move from a classical ontology (where the verb to be is defined in spacetime) to a quantum ontology (where the verb to be is defined in Hilbert space), or ii) to allow for a multi-level ontology [15], reassured by the notion of classical real patterns emerging from the underlying quantum world [18].
@@ 4.2 What are the possibilities?
Another important aspect of this discussion is the expected plurality of entities. If we argue for the universe as a whole, either classically or quantum-mechanically, we talk about things in the singular and focus on the holistic aspects of either spacetime or Hilbert space, their global state, and Hamiltonian as what ultimately exists. The partition of these spaces into what we typically see as classical and quantum subsystems is viewed as emergent, despite the persistence of real patterns associated with emergent entities [19, 18]. From this perspective, the second part of the ontological project becomes superfluous, as there is no meaningful notion of relations for a set composed of a single element, except the trivial comparison with the empty set [^20] An instance of this comparison in physics can lead to questions such as “Why is there something rather than nothing?” [30].. A recent defense of this position can be found in [31]. This holistic approach will be contrasted with the atomic one, where there exists more than one single irreducible entity and relationships among them.
@@ 4.4 Quantum Ontology
The context of quantum mechanics makes everything much more interesting. In quantum mechanics, the primary elements of reality are not particles localized in spacetime but quantum states described by vectors in Hilbert space. These states can exist in superpositions and exhibit spatiotemporal non-local correlations, challenging the classical notion of independence. Adopting a quantum ontology thus requires rethinking the independence and interaction of parts. This perspective forces a reconsideration of how we define and measure the components of a system, as well as how we understand their interactions, topics we will explore in Sec. 7 and Sec. 8. Quantum mechanics seems to suggest a more holistic view, where the properties of the whole system cannot always be decomposed into the properties of its parts. This reasoning also justifies approaches to quantum cosmology and views where reality is seen as an evolving vector in Hilbert space [31].
@@
From this perspective, the main reason to consider this ontology is to attribute a sense of reality to entities that satisfy the Schrödinger equation, in contrast to quantum subsystems that seemingly demand some form of non-unitary evolution, apparent or not. Thus, the measurement problem will play a crucial role in how we approach this topic in Sec. 8.
@@
explores the possibility of maintaining a form of atomism within the framework of quantum mechanics, where the universe is composed of distinct quantum subsystems. Unlike classical atomism, which relies on spatial localization and independence, each quantum subsystem is associated either with factors in Hilbert space for finite-dimensional systems (Sec. 2) or by commuting local operator algebras in algebraic quantum field theory (Sec. 5). However, as we will discuss in Sec. 6.2, quantum gravity seems to undermine these decompositions, at least as long as they are understood as invariant under the unitary evolution of the global Hilbert space to which they belong. We will revisit this in Sec. 8, where we will offer an alternative way to bridge holistic and atomistic quantum ontology, providing a new perspective on the measurement problem.
@@ 4.5 How do we decide?
As mentioned briefly above, we seem to be trapped in spacetime with classical measurements. Thus, given our invocation of scientific realism, which employs rationales such as the no-miracles argument [32] to attribute a sense of reality to theoretical terms in our theories, we soon are led into problems, such as our mysterious box in Sec. 1. There, we saw that under our limited measurements, only the energy displayed by the box, we could not infer the nature of the quantum system inside the box.
As discussed in Sec. 2, atomistic classical and quantum ontologies are intrinsically interconnected by the requirement that local measurements can be performed independently, establishing a direct connection between macro- and microcausality. Now, if there is ambiguity at the quantum level about how the Hilbert space is partioned and how local the quantum subsystems are interacting, we could imagine that further criteria should be imposed to select a preferred factorization of the Hilbert space, or assignment of local algebras. This project is pursued to some extent [^22] Other ideas for selecting a preferred factorization of the Hilbert space have been recently pursued in [7, 8, e.g.], with general discussions about decompositions of Hilbert space found in [37]. in [6] where they argue that imposing a quantum mereology resulting in the most possible quantum-local Hamiltonian is (almost always) sufficient to select a particular TPS, which then leads to a specific classical limit. From the perspective of classical ontology, we can invert this picture and view their program (which could be thought of as establishing the grounds for discussing quantum mereology from a Hilbert space fundamentalist perspective) as recognizing that we indeed live in a specific classical limit and then finding the necessary criteria at the quantum level aligned with this classical limit.
@@ 5 Status-of-art, or crash course on algebraic quantum field theory
At the end of the day, we are seemingly trapped in spacetime. By this, we mean that the real patterns [19] surrounding us are best described by spacetime physics, which includes all the possible measurements we make [38, 39]. In fact, despite their complexity, measurements still amount to simply counting—such as the number of clicks in a detector or spots on a screen—from which we can infer the state of our physical systems.
@@ 5.1 What could we have hoped for?
We operate under the belief that the world is fundamentally quantum-mechanical and that the entire spacetime is associated with an underlying global Hilbert space. Here we see the relevance and direct connection between the holistic ontologies discussed in Sec. 4.2. A crucial aspect of our activities in spacetime is the idea that it can be divided into independent parts. This concept underlies the separation between a system to be measured, our devices, and even ourselves as observers. The question then becomes what is needed to ensure that the whole can be broken into independent pieces, and to what extent this can be realized in Hilbert space.
Unexpectedly, this is a deep philosophical and physical problem that also appears in many other disciplines. We do not intend to provide an exhaustive discussion of the matter, but we will highlight at least one criterion that we already met that is needed classically to define any reasonable notion of independence: causal independence. If we know that two systems cannot exchange any information in a way that all possible observers would agree on, this is a strong criterion for considering them as independent entities. Special and general relativity provide the framework in which this can be formulated. We will refer to this causal independence as spatiotemporal locality, as it establishes the condition under which we can securely say that something is independently located relative to something else (if they were not causally disconnected, there would be observers for whom the systems could be arbitrarily close to each other).
Let’s be more precise. What we have in mind are two small regions of spacetime, $U$ and $V$, which are spacelike separated and define local Cauchy surfaces. This means that there is a spacelike surface inside each region that contains all the necessary information to describe anything inside those regions.
@@
If we recall our initial mysterious box, perhaps this should not come as a surprise. Although quite often locality in Hilbert space, in the sense of local subsystems with commuting algebras, is a direct consequence of spatiotemporal locality (as discussed in Sec. 2), this is mostly our spacetime parochialism. As emphasized in recent works [40, 6, 7, 8, e.g.], the notion of spatiotemporal locality should be seen as emergent from the quantum formalism. So what can be done instead?
@@ 5.2 The Algebraic Approach
By analogy with coarse-grained emergent descriptions discussed in Sec. 3.2 and 3.3, in the space of quantum models aimed at reproducing classical spacetime measurements, is there a notion of an ‘ensemble’ of models that all converge to the same set of numbers one obtains by taking measurements in the lab? Is there a notion of model coarse-graining that needs to be considered because, similar to our accessible macrostates in thermodynamics, we only have access to classical measurements in the end? We will return to these questions in Sec. 8, but we have already seen that in the case of dualities, this is exactly the situation: different sets of commuting subalgebras can be used to arrive at the same classical measurements, even though they give rise to different spatiotemporal local structures, as observed in Sec. 2.
This highlights the subtleties involved in defining quantum subsystems and the emergent nature of spatiotemporal locality from a more fundamental quantum framework. And now, at least we finally understand why the maps between the different sets of spin sites were non-locally related, since the choice of commuting subalgebras is one of the criteria that leads to the notion of independent subsystems.
@@ 5.4 More ambiguities!
If we could turn off the gauge interaction, or effectively deactivate the scalar field charge, we could consider the complex scalar field components as a local observables (assuming some smearing) associated with different spacetime regions, such that for any two spacelike-separated regions $x\in U$ and $y\in V$, microcausality holds:
@@
where $\Gamma$ is a curve connecting $x$ and $x^{\prime}$ and it represents a Wilson line connecting two field excitations of opposite charges. Then, as long as $D(x,x^{\prime})$ can be localized in $U$, it will commute with analogous operators defined in other spacelike-separated regions, and once again we find a gauge-invariant algebra of observables that is microcausal.
@@ 6 What about gravity?
Any good story starts with a mystery and, at some point, introduces a villain. Our understanding of gravity has dramatically changed since Newton’s theory. Initially, space and time were deemed absolute, and gravity was yet another force acting on this absolute stage. However, we have since learned with Einstein that the stage itself is dynamical, changing depending on the matter-energy act at play, and this very dynamics is responsible for gravitational phenomena.
@@
Thus, it is quite clear that Einstein’s perspective does not take space, and perhaps time, as independent substances. However, Einstein’s general relativity did not quite perfectly achieve this strict version of relationalism, as it does allow vacuum solutions, but it was a significant step in that direction. Regardless, general relativity is extremely successful and underlies both astrophysics and modern cosmology, each presenting a large body of evidence towards its validity across different physical regimes.
@@
Nonetheless, many lessons have been learned so far. But we will focus on one lesson that will lead us to reconsider several approaches to important problems in different areas of physics in Sec. 7. This is the problem of defining local gauge-invariant observables in linearized quantum gravity.
@@ 6.1 Operational Measurement Limitation
Let’s start intuitively [52, see interview with Nima Arkani-Hamed]. Imagine you are back in your lab, and you want to make a perfectly sharp measurement, or at least as sharp as possible. The problem is how to store that information in your apparatus. Quantum mechanically, you could imagine having an infinite chain of qubits to record any measurement using binary digits. This would take a lot of space, so perhaps you have access to a quantum field that has an infinite-dimensional Hilbert space at each point in space. Either way, you need an infinite-dimensional Hilbert space for a perfectly sharp measurement. While this may not be feasible in practice, there is nothing in quantum mechanics that prevents you from imagining it in principle.
@@
Now, let’s explore more concretely what happens when we try to define local gauge-invariant observables in gravity.